### v3

# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

Abstract
Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery. Mathematical competitions, in particular, present two challenging problem types: theorem-proving, which requires rigorous proofs of stated conclusions, and answer-construction, which involves hypothesizing and formally verifying mathematical objects. Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation. We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving. We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations. On the ConstructiveBench dataset, ECP improves the accuracy of answer construction from a Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with the gpt-4.1-mini model. Moreover, combined with ECP’s constructed answers, the state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines. Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
1University of Toronto  2Vector Institute  3Georgia Institute of Technology  4Google DeepMind
1 Introduction
Mathematical reasoning is fundamental to artificial intelligence (Newell and Simon, 1956), enabling significant progress in domains such as mathematics education, formally verified software, and mathematical research itself (Li et al., 2024). Recent advancements show that LLMs, when guided by careful prompting and tool integration, excel in solving straightforward high-school mathematics problems, but substantial challenges remain in problems requiring formal proof or the construction and verification of complex mathematical objects (Yang et al., 2024). MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer constructions.
Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
Answer-Construction Problem:
Find all functions $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ such that
$\forall x,y\in\mathbb{R}^{+}$, $x\bigl{(}f(x)+f(y)\bigr{)}\geq\bigl{(}f(f(x))+y\bigr{)}f(y)$ Answer: All $f(x)=\frac{c}{x}$ for some $c>0$. Proof:
Theorem-Proving Problem:
Let $x_{1},\dots,x_{2023}$ be distinct positive numbers such that for all $n\in\{1,\dots,2023\}$,
$\sqrt{\left(\frac{1}{x_{1}}+\dots+\frac{1}{x_{n}}\right)\,(x_{1}+\dots+x_{n})}\in\mathbb{Z}$ Prove that $x_{2023}\geq 2024$. Proof:
Table 1: Examples of competition-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021). While LLMs are effective at creative and exploratory reasoning, they frequently generate unverifiable and incorrect assertions due to hallucinations (Huang et al., 2025). Symbolic reasoning tools, such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008)), serve as the logical backbone of automated reasoning tools but cannot feasibly discover closed-form answers in the vast space of candidates.
Inspired by mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer-construction problems in math competitions; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
2 Related Work
Formal Theorem Proving
Formal methods such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008), CVC5 (Barbosa et al., 2022)) and first-order provers (e.g., Vampire (Kovács and Voronkov, 2013), E (Schulz, 2002)) are widely used for formal verification and logical reasoning. Interactive theorem provers (ITPs) like Isabelle (Paulson, 1994) and Lean (Moura and Ullrich, 2021) allow step-by-step formal proof development, often integrating automation via tools such as Sledgehammer (Böhme and Nipkow, 2010) and LeanCopilot (Song et al., 2024).
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench. Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
LLMs for Mathematical Reasoning
Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs. Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity. Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems. Reinforcement learning-based approaches (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals. Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools. SATLM (Ye et al., 2023) extends this framework by using a Z3 symbolic solver to validate logic-based task specifications.
Evaluation Benchmarks
Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format. Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision. In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation. Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light. LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalizes the Mathlib4 corpus back to natural language.
3 Methods
3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$. Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
In contrast to theorem-proving tasks, an answer-construction task asks for constructing mathematical objects that satisfy specified constraints, without an explicitly stated conclusion. We define an answer-construction task by the following components:
Context Variables:
The unassigned context variables $a:=(a_{1},\dots,a_{k})$ (with type $\alpha:=\alpha_{1}\times\dots\times\alpha_{k}$) in the problem.
Context Property:
The predicate $P:\alpha\rightarrow\text{Prop}$ that the context variables satisfy.
Answer Variables:
The unknown answer $b$ (with type $\beta$) to be constructed.
Answer Function:
$f:\alpha\to\text{Set }\beta$ mapping each context instance to a set of valid answers.
Problem Predicate:
A Boolean predicate $Q:\alpha\times\beta\to\text{Prop}$ such that $P(a,b)=\text{True}$ if and only if $b$ is a valid answer under context $a$.
We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$. The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated as follows in Lean’s dependent type theory:
$\exists f:\alpha\to\mathrm{Set}\ \beta,\ \forall(a:\alpha),\ P(a)\rightarrow\forall(b:\beta),\ Q(a,b)\leftrightarrow b\in f(a).$
[…]
$\exists f:\alpha\rightarrow\beta,\ \forall(a:\alpha),\ P(a)\rightarrow\bigl{(}\forall(b:\beta),\,Q(a,b)\leftrightarrow b=f(a)\bigr{)}.$
When there are no context variables, the formulation reduces to $\exists f:\text{Set }\beta,\ \forall b:\beta,\ P(b)\leftrightarrow b\in f.$
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in a closed-form, algebraically simplified representation that generalizes across instances. Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g., setting $f(a,b)=\{x\mid ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}\mid x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1. Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover. We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification. Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
3.2 Autoformalization Pipeline
Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean. Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail to produce precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, as illustrated in Figure 1. The pipeline comprises three core components: (1) autoformalizations with LLMs, (2) Lean knowledge retrieval, and (3) error-guided feedback loops.
Figure 1: Overview of the autoformalization pipeline with error-based feedback and documentation retrieval: An informal statement is first processed by the LLM to produce a formalization. If the result fails either the Lean compilation check or the semantic check by an LLM judge, error messages and relevant Lean documentation are retrieved and fed back to the LLM. The process continues iteratively until the LLM produces a formalization that passes both checks.
Autoformalizations with LLM.
We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini and deepseek-v3, for autoformalization. To better align these models with formalizations in high-school and undergraduate mathematics, we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023). Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
Lean Knowledge Retrieval.
LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated or scarce training data in formalized statements. To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and the Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023). After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries. Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024). When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances formalization accuracy and reduces hallucination rates.
Error-Guided Feedback Loop.
Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies. Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations. This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers. Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, ensuring consistency across the formalization process. Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
3.3 Enumerate–Conjecture–Prove Framework
ECP systematically integrates LLMs’ capabilities for coding and mathematical reasoning with formal theorem proving within the Lean proof assistant. Figure 2 visualizes the workflow of ECP applied to an example problem.
Figure 2: Illustration of the ECP framework applied to a Balkan MO Shortlist problem in ConstructiveBench.
Enumerate Stage.
In the initial Enumerate stage, the “enumerator" LLM generates executable Python scripts designed to enumerate explicit examples or special cases of the given mathematical conditions, using the prompts in Appendix B. These enumeration scripts systematically evaluate finite subsets of context variables under computational constraints (timeout of 60 seconds and a maximum of 100 candidate answers) to gather concrete data points and identify initial numerical or structural patterns.
Conjecture Stage.
Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern-finding in enumerated answers. These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification. We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
Prove Stage.
In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task. The theorem is then verified using specialized LLMs fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025). Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
  1. 1.
    simp: Simplifies expressions by rewriting.
  2. 2.
    aesop (Limperg and From, 2023): Proves the goal via a best-first search proof engine.
  3. 3.
    nlinarith: Solves nonlinear arithmetic problems over the integers and rationals.
  4. 4.
    ring: Proves equalities in commutative (semi)-rings.
  5. 5.
    norm_num: Evaluates and proves numeral equalities and inequalities.
Successful verification by Lean concludes the ECP process.
4 ConstructiveBench Dataset
In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination.
4.1 Data Sources and Scope
ConstructiveBench is curated by aggregating and rigorously filtering problems from reputable mathematical sources. Primary sources include official competition archives such as AMC 12 A/B (from 2000 onward), AIME I/II, HMMT, regional Olympiads, and the IMO Shortlist and Longlist. Additionally, we integrate data from established informal math datasets including OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024).
The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions. It covers diverse mathematical domains including algebra, number theory, combinatorics, probability, and calculus. Rather than explicit domain labeling, we employ structural filtering to retain exclusively answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
((a)) Problem sources
((b)) Problem domains
Figure 3: Problem sources and problem domains in ConstructiveBench. Only the top 11 problem sources and top 7 problem domains are shown; the remaining are grouped under ‘Other’.
4.2 Data Deduplication
The dataset initially aggregated 5,555 problems. To remove duplication, we utilized both string-based matching and embedding-based duplicate filtering methods, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold. This process resulted in 5,316 unique and diverse problems.
4.3 Problem Autoformalization
We used our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into a formally stated Lean theorem. Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with the gpt-4o model serving as an LLM judge to verify semantic consistency.
Each problem underwent up to $T=5$ refinement iterations per model. This procedure yielded a well-formed subset of 3,431 problems, whose formalizations compile successfully in Lean, pass the LLM-based semantic check, and are formatted according to the template. Each entry in ConstructiveBench includes the informal and formal problem statements, the ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty. Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench. The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various types of mathematical problems.
Furthermore, to address potential data contamination—where LLMs might inadvertently have memorized training examples—we explicitly include a carefully selected test subset of 92 problems drawn from formalized data created after June 2024, which is beyond the knowledge cutoff date of most LLMs (including those we will evaluate: gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
Figure 4: Left: Distribution of answer types in ConstructiveBench. Only the top 8 types are shown; the remaining types are grouped under ‘Other’. Right: Example dataset entry from ConstructiveBench.
5 Empirical Evaluation
The objective of our empirical evaluation was to answer the following questions.
  1. 1.
    Autoformalization Quality: Can ECP efficiently produce error-free and semantically faithful formalizations?
  2. 2.
    Answer-Construction Accuracy: Can ECP enable more accurate answer construction compared to CoT baselines?
  3. 3.
    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success rates on answer-construction problems compared to purely symbolic or LLM-only approaches?
5.1 Experimental Setup
Autoformalization.
To evaluate the quality and efficiency of our autoformalization pipeline, we measure the percentage of problems that successfully compile in Lean and are validated by our LLM judge (gpt-4o). We sequentially run up to $T=5$ iterations for three models: fine-tuned gpt-4o-mini, deepseek-math-7b-finetune, and deepseek-v3.
Answer-Construction.
For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers). We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, as well as deepseek-v3. The models are given the formalized Lean statement and the type of answer to be constructed. We compare the ECP framework against a baseline Chain-of-Thought (CoT) method (i.e., without enumeration and pattern finding) for answer construction. The constructed answer is considered correct if and only if the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics: simp, aesop, nlinarith, ring, norm_num).
Theorem-Proving.
We evaluate the theorem-proving tasks using state-of-the-art provers, comparing baseline symbolic methods with specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B. Evaluations are performed on one A100 GPU and 64 CPU cores for Lean verification, using the hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048, and the Pass@32 metric. The experiment includes two conjecturing setups (CoT or ECP), three different models for conjecturing, and three different prover models for proving, resulting in a total of 18 experiments. We use 64 AMD EPYC 7453 CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours on an A100 80 GB GPU (2 hours per setup).
5.2 Results
Autoformalization.
Out of the 5,316 problems in ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks, yielding an overall success rate of 64.6%. The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
Answer-Construction.
We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified ConstructiveBench problems and on the subset of 337 answer-construction problems from the 640 problems in PutnamBench. Answer correctness was rigorously verified through formal proofs in Lean.
The results, summarized in Table 2, clearly demonstrate ECP’s advantage over standard LLM-only methods. On the ConstructiveBench dataset, ECP substantially improved answer accuracy for gpt-4.1-mini from 14.54% (baseline) to 45.06%, gpt-4.1-nano from 17.02% to 32.59%, and deepseek-v3 from 19.06% to 43.81%. Notably, gpt-4.1-mini achieved the highest accuracy within the ECP framework.
On the uncontaminated subset of 92 problems created after June 2024, performance improvements were also pronounced: gpt-4.1-mini improved from 5.43% to 31.52%, gpt-4.1-nano from 5.43% to 20.65%, and deepseek-v3 from 4.35% to 29.35%. These significant gains demonstrate ECP’s robust generalization capabilities, particularly for novel problems beyond the training cutoff.
Model / Dataset ConstructiveBench ConstructiveBench (After Cutoff) PutnamBench
gpt-4.1-nano 17.02% / 32.59% 5.43% / 20.65% 7.72% / 11.57%
gpt-4.1-mini 14.54% / 45.06% 5.43% / 31.52% 15.13% / 23.15%
deepseek-v3 19.06% / 43.81% 4.35% / 29.35% 18.99% / 31.75%
Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and on 337 PutnamBench answer-construction problems.
Theorem-Proving.
We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem proving using the same ConstructiveBench problems. Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
Prover Model
Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimina-Prover-7B-Preview DeepSeek-Prover-V2-7B
gpt-4.1-nano CoT 7.41% 10.46% 11.45% 15.42%
ECP 8.11% 13.61% 16.03% 20.61%
gpt-4.1-mini CoT 6.42% 8.07% 9.24% 13.00%
ECP 9.86% 15.83% 18.51% 25.01%
deepseek-v3 CoT 6.04% 9.62% 10.49% 15.33%
ECP 9.13% 20.26% 18.27% 24.63%
Table 3: Theorem-proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
The summarized results on the ConstructiveBench problems are presented in Table 3. Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini. Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B. Additionally, on the subset of 337 answer-construction problems from PutnamBench, using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, compared to 0.59% with CoT for answer construction. Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model–prover combinations.
6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. Our framework and dataset depend on the quality of autoformalization produced by black-box LLMs. In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity, it is time-consuming and difficult to scale. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem. Future work should explore improved evaluation metrics for formalization quality and expand datasets that better align informal and formal representations of mathematical problems.
Theorem-Proving. Our proving stage uses off-the-shelf theorem provers, which leaves significant room for optimization. In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space. Fine-tuning LLMs or developing search-based algorithms for the formal answer-construction task thus represents a promising avenue for next-generation mathematical AI models.
7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedback. In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%. Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer-construction mathematics problems. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.
Appendix
Appendix A Lean Definition Filtering
To filter out the rarely-used definitions in autoformalization, we prioritize the definitions and theorems within the following namespaces in Lean:
Nat, Int, Rat, Real, Complex, ENat, NNReal, EReal, Monoid, CommMonoid,
Group, CommGroup, Ring, CommRing, Field, Algebra, Module, Set,  Finset,
Fintype, Multiset, List, Fin, BigOperators, Filter, Polynomial, Order,
SimpleGraph, Equiv, Embedding, Injective, Surjective, Bijective, Topology
Appendix B Prompt
B.1 Prompt for Formalization
[…]
‘‘‘lean
your code
B.2 Prompt for Enumerator Coding
LLM Prompt I have a difficult high-school competition-level math problem. Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer. Use the following format for the delimited output:
‘‘‘python
your code
B.3 Prompt for Conjecturing
LLM Prompt I have a difficult high-school competition-level math problem in Lean, which asks to find a answer set with a mathematical proof. Your task is to clearly state the final answer in Lean exprssion with explanations.You do not have to derive the whole proof. I have enumerated the answer(s) via program as a hint for you, and you may look at the pattern in it to help propose the final answer. Your final answer should be in closed form in Lean and generalized from enumerated answer that can eliminate the existential quantifier. You must not echo the problem statement in your answer. Use the following format for the delimited output:
‘‘‘lean
your code